a student holds two weights in his hands while spinning on a rotating chair. what happens to his rotation rate when he drops the weights?

increases
decreases
remains the same
Answer #1

Total angular momentum of student holding weights, J = I*w.

here w = rate of rotation

I = total moment of inertia

I is proportional to total mass(I=m*r*r) ; m=mass , r= distance of mass from axis of rotation.

When weights are dropped I decreases, so to conserve the angular momentum(keeping the value I*w fixed), w must increase.